In a genetics experiment on​ peas, one sample of offspring contained 434 green peas and 602 yellow peas. Based on those​ results, estimate the probability of getting an offspring pea that is green. Is the result reasonably close to the value of 3/4 that was​ expected?

The probability of getting a green pea is approximately?

​(Type an integer or decimal rounded to three decimal places as​ needed.)

Is this probability reasonably close to 3/4​?

Choose the correct answer below.

A.

No, it is not reasonably close.

B. Yes​, it is reasonably close.
Answer #1

Probability of getting a green pea = 434/(434+602)

= 0.4189

H0: p = 3/4 = 0.75

H1: p 0.75

= 0.419

n = 434+602

= 1036

Test-statistics:

p-value ~ 0 < 0.05 (significance level)

i.e. we can reject the null hypothesis and can say that probability of getting an offspring pea that is green is different than the value 0.75.

So the correct answer is:

A. No, it is not reasonably close.

Please upvote if you have liked my answer, would be of great help. Thank you.